A 10% increase in the price of pizza causes a 10% drop in the quantity of both pizza and beer sold. Describe elasticities and the nature of the two products.
Answer #1

Answer

elasticity of demand =%change in quantity of the good /%change in the price of the good

elasticity of demand for Pizza=-10/10=-1=1 (absolute value)

the demand is unit elastic

======

Cross price elasticity of demand =%change in quantity of one good/ %change in the price of other good

=-10/10

=-1

the cross-price elasticity of demand is negative so the goods are complement goods.